OpenStudy (jkbo):

Please help.Without graphing, classify the following system as independent, dependent, or inconsistent. -x-2y=-7 5x-y=2 a.independent b.dependent c.inconsistent

OpenStudy (anonymous):

do u know matrices?

OpenStudy (jkbo):

@chris00 no who is ?

OpenStudy (anonymous):

ok, nevermind that. i'll give u a head start. if the system has solutions then it is independent. if the system has infinite solutions, then it is dependent if the system has no solutions it is inconsistent

OpenStudy (jkbo):

It would be dependent, right? thats what I just got

OpenStudy (anonymous):

how did you get that?

OpenStudy (jkbo):

when I solved it , it had infinite solutions. I know its not inconsistent

OpenStudy (anonymous):

tell me how you solved this

OpenStudy (jkbo):

I subtracted two from each side,

OpenStudy (anonymous):

lol wat? we are trying to find solutions for x and y that satisfy both equations

OpenStudy (jkbo):

O ok well I'm not sure how to solve this then

OpenStudy (anonymous):

if we isolate x on one side we get x=7-2y substituting this into the second equation, 5(7-2y)-y=2 35-10y-y=2 -11y+35=2 11y=33 y=3 so using y=3 sub this into x=7-2y x=7-2(3) =7-6=1 (1,3) is a solution for this system, so what does it tell us about the nature of the system?

OpenStudy (jkbo):

its independent because it has solution

OpenStudy (anonymous):

correct

OpenStudy (jkbo):

thank you